if f = 4i – 3j then find prependicular vector to f and solve step by step
 

Grade:12th pass

3 Answers

Saurabh Koranglekar
4 years ago
Dear student

Let q be another vector ai +bj

f.q = 0

4a- 3b = 0 is condition

Regards
Arun
4 years ago

Talking about two perpendicular vectors,

The dot productof both is zero. Always.

a•b= ab cosα

So, the vector is 4i-3j

By the above relation we get,

(4i-3j).

(4i-3j).(ai+bj)=0

4a-3b=0

4a=3b

a/b=3/4

Therefore the vector perpendicular to given vector is 3i+4j.

aswanth nayak
4 years ago
Dear student,
Let x be a vector ai+bj
The dot product of both the vectors is ZERO

Talking about two perpendicular vectors,

a•b= ab cosα

So, the vector is 4i-3j

By the above relation we get,

(4i-3j).

(4i-3j).(ai+bj)=0

4a-3b=0

4a=3b

a/b=3/4

Therefore the vector perpendicular to given vector is 3i+4j.

 

hope this helps you

Regards